OpenStudy (anonymous):

I WILL FAN AND MEDAL y=4x-18 rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

Re-wrtie it so that the x and the y are on the same side.

OpenStudy (anonymous):

4x+y=18?

OpenStudy (anonymous):

How did you do that ? it looks like you lost a negative (or minus) sign.

OpenStudy (anonymous):

oops i meant to put 4x+y=-18

OpenStudy (anonymous):

Ok now that is a little better ... but how did you get the 4x over to the right side

OpenStudy (anonymous):

sorry how did you get 4x over to the LEFT side?

OpenStudy (anonymous):

-4x+y=-18

OpenStudy (anonymous):

I forgot I had to subtract haha

OpenStudy (anonymous):

There you go... now this all depends on your book/course/teacher ... some consider Standard form Ax+By=C and some consider it Ax+By+C=0 ... (in all cases A,B,C must be integers - not fractions or decimals) ...do you know which your book/course/teacher wants?

OpenStudy (anonymous):

Ax+By=C or Ax+By+C=0?

OpenStudy (anonymous):

Ax+By=c

OpenStudy (anonymous):

Ok so you are almost there. Usually the A is supposed to be greater than or equal to 0. But you have it as -4 so when this happens just multiply both sides of the equation by -1 and that solves that problem and the equation looks different but is really the same. so go ahead and fix your equation.